OpenStudy (anonymous):

An angle measures 30 degrees more than twice its complemet. Find the measure and its complement. Help?

OpenStudy (anonymous):

This is what I got. Don't know how thought sorry if I get it wrong.. Let x = the angle Then x -30 = the measure of the angle's complement Since they are complementary angles, the angle + the angle's complement = 90 degrees Thus x + (x-30) = 90 x + x - 30 = 90 2x - 30 = 90 2x = 120 x = 60 degrees for the angle x-30 = 60 -30 = 30 degrees for the angle's complement.. So yea

OpenStudy (anonymous):

Shouldn't it be plus 30

OpenStudy (anonymous):

Do you have answer choices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok.. Hold on

OpenStudy (anonymous):

I'm stumped.

OpenStudy (anonymous):

You can try +30 and see if it is correct.

OpenStudy (anonymous):

Lol ok thanks for your help